OpenStudy (anonymous):

If 6xa=54 and b-a= 14, what is A x B?

MsBrains (ms-brains):

54 * 14 is what you are asking, right?

OpenStudy (anonymous):

6* 9= 54 and 19-9=14

OpenStudy (anonymous):

so what is 9*19?

OpenStudy (anonymous):

What is a x b then

OpenStudy (anonymous):

171

OpenStudy (anonymous):

\[6a=54 <=> a=9\] Insert value for a: \[b-9=14 <=>b=23\] \[b*a=9*23=207\]

OpenStudy (md152727):

Right!

OpenStudy (anonymous):

I'll use a calculator just to check

OpenStudy (retireed):

When does 19 - 9 = 14 ????

OpenStudy (anonymous):

woops! meant 207

OpenStudy (anonymous):

23-9

OpenStudy (anonymous):

Thanks everyone